Assume the Malaysian Ringgit, MYR exchange rate decreased from ฿7.48 Thai Baht in November 2019 to ฿7.00 Thai Baht in November 2020, and has experience an increase from New Taiwan Dollar, NT$6.50 in November 2019 to NT$6.98 in November 2020.

a) What was the value of the New Taiwan Dollar, NT$ in terms of Malaysian Ringgit, MYR in November 2019 and in November 2020? Did the NT$ appreciate or depreciate against the MYR over the year from November 2019 to November 2020?

b) What was the value of the Thai Baht in terms of Malaysian Ringgit, MYR in November 2019 and in November 2020? Did the Thai Baht appreciate or depreciate against the MYR over the year from November 2019 to November 2020?
1
Expert's answer
2020-12-25T15:25:45-0500

a)

November, 2019


"\frac {1}{6.5}=0.154"

November, 2020


"\frac {1}{6.98}=0.143"

Value is lower.


b)

November, 2019


"\frac{1}{7.48}=0.134"

November, 2020


"\frac {1}{7.00}=0.143"

Value is higher